OpenStudy (anonymous):

determine the value for x and y x+y=6 and 2x-2y=16

OpenStudy (anonymous):

x+y=6 and 2(x-y) = 16/2 x=7 y=-1

OpenStudy (anonymous):

y=6-x 2x-2(6-x)=16 2x-12+2x=16 4x-12=16 4x=16+12 4x=28 x=28/4=7 y=6-7=-1